# ScVLM: a Vision-Language Model for Driving Safety Critical Event Understanding

### v1

@@ Abstract
Accurately identifying, understanding, and describing driving safety-critical events (SCEs), including crashes and near-crashes, is crucial for traffic safety, automated driving systems, and advanced driver assistance systems research and application. As SCEs are rare events, most general Vision-Language Models (VLMs) have not been trained sufficiently to link SCE videos and narratives, which could lead to hallucination and missing key safety characteristics. To tackle these challenges, we propose ScVLM, a hybrid approach that combines supervised learning and contrastive learning to improve driving video understanding and event description rationality for VLMs. The proposed approach is trained on and evaluated by more than 8,600 SCEs from the Second Strategic Highway Research Program Naturalistic Driving Study dataset, the largest publicly accessible driving dataset with videos and SCE annotations. The results demonstrate the superiority of the proposed approach in generating contextually accurate event descriptions and mitigate hallucinations from VLMs.
@@ 1 Introduction
In the domain of autonomous driving, VLMs have demonstrated strong and robust capabilities in perception, scene understanding, decision-making, and adaptability to novel scenarios [1, 2, 3, 4]. Related to driving safety, VLMs proficiently interpret environmental information surrounding the vehicle and possess foundational insights into traffic accidents and potential risk factors [2, 4]. Despite these advances, VLMs still face challenges in accurately identifying safety-critical events (SCEs), including crashes and near-crashes. Furthermore, understanding the nature of these SCEs, such as conflicts with a lead vehicle, remains elusive. This information is crucial for assessing driving safety.
Figure 1: Example scene understanding result by VideoLLaMA2
Figure 1 illustrates the capabilities of the advanced VLM, Video Large Language Model for Learning from All Modalities 2 (VideoLLAMA2) [5], in understanding SCEs. This model exhibits a excellent understanding of static environmental contexts, including weather conditions and the immediate surroundings. However, its ability to discern dynamic elements crucial for SCE analysis—such as distinguishing between crash and normal driving scenarios or identifying the nature of conflicts (e.g., with a leading or parked vehicle)—is still constrained. These findings underscore the necessity for improved models capable of more effectively interpreting dynamic information in SCE videos.
The limitations observed in VideoLLAMA2’s performance on SCE analysis can be attributed to two key factors. The rarity of SCEs in real-world scenarios results in insufficient training data for general VLMs to establish connections between SCE videos and corresponding narratives [6, 7]. The scarcity of relevant training examples can lead to hallucination and the omission of crucial safety characteristics in the model’s interpretations. Additionally, the abstract nature of event types and conflict types poses a significant challenge for VLMs to accurately identify [8].
This work introduces a novel hybrid approach that utilizes supervised learning, contrastive learning, and VLM for driving video understanding, with emphasis on SCEs. For event type identification, including crashes, tire strikes, near-crashes, and normal driving, supervised learning techniques are used due to their task-specific effectiveness. For conflict type identification, contrastive learning is employed, taking advantage of semantic label dependency on labels with rich text information. To interpret the surrounding environmental context, VLM is utilized for its capability to accurately identify concrete objects in the video. Finally, VLM combines the predictions from the supervised and contrastive learning approaches with the environmental context to generate a coherent event description in fluent language.
@@
The evaluation of the proposed approach utilized data from the Second Strategic Highway Research Program (SHRP 2) Naturalistic Driving Study (NDS), which is the largest publicly accessible NDS dataset to date, containing over 1,000,000 hours of continuous driving data [9]. The SHRP 2 NDS data includes rich driving information from multiple cameras, kinematic sensors, radar, and GPS. From the continuous driving data, a dedicated project was conducted to identify SCEs and randomly selected normal driving baselines [9], including four distinct event types: crashes, tire strikes, near-crashes, and normal driving baselines. SCEs went through a rigorous data annotation process to extract the nature of conflict. The annotation provides detailed conflict type labels for SCEs, covering scenarios like conflicts with a lead vehicle, single-vehicle conflicts, and conflicts with a vehicle turning into another’s path in the same direction. This rich dataset is ideal for evaluating the effectiveness of the proposed hybrid approach.
@@ 2 Related Works
VLM for Driving Scene Understanding VLMs combine visual and language processing to interpret driving scenarios and aid decision-making. DriveVLM, which incorporates reasoning modules for scene description and analysis, addressing spatial reasoning and computational challenges by proposing a hybrid system combining VLMs with traditional autonomous driving pipelines [3]. DriveScenify, utilizing advanced VLMs to generate contextually relevant responses based on driving scene videos, aiming to enhance urban mobility and road safety [10]. Shoman et al. [2] propose a parallel architecture that integrates object detection, tracking, and natural language generation to produce detailed descriptions of traffic events, thereby improving traffic safety through comprehensive event analysis. Jain et al. [11] integrate VLMs with multi-sensor data to enhance the comprehension of traffic dynamics and interactions among road users and infrastructure.
While there has been a significant expansion in research on VLMs for general scene understanding in driving contexts, the specific focus on SCEs, which are crucial for enhancing safety and reliability in autonomous vehicles, is relatively under-explored. Some researches that mention crashes or traffic accidents [2, 4, 3] do not explore the intricacies of these events in depth.
Supervised Learning and Contrastive Learning Supervised learning and multi-modal contrastive learning are two popular approaches for driving video scene classification tasks [12, 13, 14, 15, 16]. Supervised learning relies on one-hot or figure-coded labels to train models [17], while multi-modal contrastive learning, particularly in a video-text manner, leverages the relationships between different modals of the data to learn useful representations [18]. In supervised learning, state-of-the-art and efficient methods such as SlowFast [19], Swin Transformer [20], and TimeSformer [21] have proven effective for video scene understanding. Meanwhile, in contrastive learning, inspired by CLIP [22], notable approaches like X-CLIP [23] and ActionCLIP [24] excel in video understanding, particularly in few-shot tasks. The X-CLIP introduced a lightweight cross-frame attention mechanism and proposed a video-adaptive textual prompting scheme to handle video-text datasets [23]. The ActionCLIP introduced textual and visual adapters to enhance the model’s ability to process and understand text and video modalities [24].
@@ 3 VLM-based Driving SCE Analysis
To generate a comprehensive and accurate description of SCEs, the proposed approach is divided into three stages, as illustrated in Figure 2. The first stage employs a supervised learning approach to analyze the front-view videos and classify four types of events – crashes, tire strikes, near-crashes, and normal driving. In the second stage, a contrastive learning approach is utilized to identify 16 types of conflicts, such as conflict with a lead vehicle, a parked vehicle, or a following vehicle. The final stage integrates the event and conflict type information into a VLM to generate a comprehensive description of the event.
Figure 2: Overall Model Process
@@ 3.1 Supervised Learning for Event Type Classification
Supervised learning for event type classification from video is a 1-of-N vote problem, as illustrated in Figure 3. This type of model takes a video as input and feeds it through a video encoder to generate video representation. The representation is subsequently processed by a classifier to produce prediction scores. The model is optimized by minimizing the cross entropy loss based on the prediction scores. Formally, given an input video $x_{i}$ and a label $y_{i},i=1,2,...,N$ from a predefined set of labels $Y$, supervised learning approaches typically train a model to estimate the conditional probability $P(y|x,\theta)$.
@@
The supervised learning approach employs a video encoder $g_{V}$, which extracts representations for video data. Then the classifier projects the video representations into the space with the dimension of labels to obtain the prediction scores:
$s(x)=\text{Classifier}[g_{V}(x)]$ (1)
@@
$\begin{split}L=\text{Cross Entropy}[s(x),y]\end{split}$ (2)
@@ 3.2 Contrastive Learning for Conflict Type Classification
Contrastive learning approach is illustrated in Figure 4. This approach processes a video-text pair as input. The input video is fed into the video encoder to generate video representations. Concurrently, the label text is fed into the text encoder to obtain text representations. Contrastive learning approach computes a similarity score matrix between the video and text representations and is optimized by minimizing the loss between this similarity matrix and the ground-truth video-text pair matrix.
@@
In contrastive learning, the video classification task is redefined as predicting the probability $P[f(x,y)|\theta]$, where $y$ represents the original label texts and $f$ denotes a similarity function. Subsequently, the inference becomes a matching process, with the label texts having the highest similarity score being the outcome:
$\hat{y}=\arg\max_{y\in Y}P[f(x,y)|\theta]$ (3)
Contrastive learning approach employs separate encoders $g_{V}$ and $g_{T}$ for videos and label texts within a dual-stream framework. The video encoder $g_{V}$ extracts spatio-and-temporal representations for video data, while the language encoder $g_{T}$ captures representations from label texts. To bring matched video and label representations closer, the similarity score is defined using cosine distances:
@@
$\begin{split}p_{x\to y_{i}}(x)=\text{SoftMax}[s(x,y_{i})]\\ p_{y\to x_{i}}(y)=\text{SoftMax}[s(y,x_{i})]\end{split}$ (5)
The ground-truth similarity scores are denoted as $q_{x\to y_{i}}(x)$ and $q_{y\to x_{i}}(y)$, respectively. The negative pair has a similarity of 0, and the positive pair has a similarity of 1. The video-text contrastive loss to be optimized is defined as
@@
Model trained by contrastive learning approach can carry out inference, as illustrated in Figure 5. When presented with a testing dataset with a label set comprising $M$ labels, the initial step involves extracting the label representations, $[t_{k}],k=1,2,...,M$, using the language encoder, $g_{T}$. Subsequently, for a given testing video, its representation $v$ is obtained through the video encoder, $g_{V}$. The similarity between $v$ and each label representation $t_{k}$ is computed using Equation (4). The label assigned to the video is determined as the one with the highest similarity score with $v$.
@@
3.3 VLM for Event Narrative Generation
In this study, VLM is utilized to generate narrative descriptions based on contextual factors, such as weather conditions, geographical location, and surrounding environment. The process involves the VLM performing inference when provided with a text prompt and the result of SCE detection, enabling accurate event description.
Formally, the video representation $V_{i}$ for a given input video $x_{i}$ is obtained using a video encoder $g_{V}$:
$\begin{split}V_{i}=g_{V}(x_{i})\end{split}$ (7)
The visual representation $V_{i}$ is subsequently processed through the Spatial-Temporal Convolution (STC) connector to capture spatial-temporal dynamics. Given a text prompt $P$, along with the predicted event type $\hat{y_{et}}$ and conflict type $\hat{y_{ct}}$ (if applicable), the output response is generated from a pretrained large language model (LLM):
$\begin{split}R=\text{LLM}[\text{STC}(V_{i}),P,\hat{y_{et}},\hat{y_{ct}}]\end{split}$ (8)
@@ 4 Application and Results
Problem Setup Utilizing the SHRP 2 NDS dataset [9], this study focuses on generating accurate narratives from front-view video of SCEs. Normal driving segments are captured a few seconds before or after SCEs within the trip as a reference level. In total, the dataset includes 1,063 crashes, 774 tire strikes, 6,782 near-crashes, and 8,497 normal driving segments, with SCEs (crashes, tire strikes, and near-crashes) classified into 16 conflict types. Each event consists of 30-second front-view video.
The proposed approach aims to address three tasks: a classification task to distinguish event types, a classification task to differentiate conflict types, and a task to generate event text narrative. To the best of the authors’ knowledge, this is the only publicly available driving video dataset with labeled event and conflict types suitable for these three tasks, thereby supporting the evaluation of the proposed approach.
@@ 4.1 SHRP 2 NDS Dataset
Table 1: Comparison of different models in event type classification
X-CLIP Contrastive 0.766 0.951 0.547 0.921 0.493 0.599 0.508
Action CLIP Contrastive 0.748 0.949 0.488 0.907 0.439 0.520 0.458
SlowFast Supervised 0.721 0.928 0.467 0.927 0.437 0.469 0.423
Swin Transformer Supervised 0.719 0.927 0.432 0.889 0.411 0.450 0.420
TimeSformer Supervised 0.713 0.945 0.468 0.920 0.448 0.487 0.459
Table 2: Comparison of different models in conflict type classification
The SHRP 2 NDS represents the largest naturalistic driving study to date, collecting data from over 3,000 participants [9, 25]. Participants’ personal vehicles were equipped with an integrated data collection system featuring four cameras (front, driver’s face, over-the-shoulder, and rear views). Video data was continuously recorded at 15 frames per second (FPS) from vehicle start to shutdown, containing over 1,000,000 hours or 70 million miles of driving data.
ID Conflict type Count
1 Conflict with a lead vehicle 3165
2 Single vehicle conflict 1441
3 Conflict with vehicle turning into another
vehicle path (same direction) 377
4 Conflict with parked vehicle 173
5 Conflict with vehicle in adjacent lane 1508
6 Conflict with vehicle turning across another
vehicle path (opposite direction) 242
7 Conflict with a following vehicle 181
8 Conflict with vehicle turning into another
vehicle path (opposite direction) 316
9 Conflict with vehicle moving across another
vehicle path (through intersection) 170
10 Conflict with animal 360
11 Conflict with vehicle turning across another
vehicle path (same direction) 65
12 Conflict with merging vehicle 121
13 Conflict with pedal cyclist 64
14 Conflict with pedestrian 163
15 Conflict with obstacle/object in roadway 176
16 Conflict with oncoming traffic 78
17 Unknown 19
Table 3: Count of SCEs by conflict types
SCEs, including crashes, tire strikes, and near-crashes, were identified through a multi-step process. This involved evaluating the kinematic characteristics of all driving data and verifying SCEs via video analysis by trained data analysts [9]. A near-crash is defined as a situation requiring an evasive maneuver by any party to avoid a crash [9], while a tire strike event is linked to a road departure incident [26]. The 8,619 SCEs are categorized into 16 conflict types, as detailed in Table 3. After excluding the ’unknown’ type, the total is 8,600. The detailed definitions of these conflict types can be found on the SHRP 2 InSight website [27].
4.2 Application
@@
Data Pre-processing The temporal location of each SCE is pinpointed using the impact timestamp from the SHRP 2 database and serves as the center point of the event [27]. A temporal window encompassing 38 video frames (equivalent to 2.5 seconds) both preceding and succeeding the event was extracted, resulting in a 5-second interval of the front-view video. The normal driving segment, randomly selected from either before or after the SCE, consists of 77 video frames to match the duration of the SCE segment.

### v2

# ScVLM: Enhancing Vision-Language Model for Safety-Critical Event Understanding

@@ Abstract
Accurately identifying, understanding and describing traffic safety-critical events (SCEs), including crashes, tire strikes, and near-crashes, is crucial for advanced driver assistance systems, automated driving systems, and traffic safety. As SCEs are rare events, most general vision-language models (VLMs) have not been trained sufficiently to link SCE videos and narratives, which could lead to hallucinations and missing key safety characteristics. Here, we introduce ScVLM, a novel hybrid methodology that integrates supervised and contrastive learning techniques to classify the severity and types of SCEs, as well as to generate narrative descriptions of SCEs. This approach utilizes classification to enhance VLMs’ comprehension of driving videos and improve the rationality of event descriptions. The proposed approach is trained on and evaluated by more than 8,600 SCEs from the Second Strategic Highway Research Program Naturalistic Driving Study dataset, the largest publicly accessible driving dataset with videos and SCE annotations. The results demonstrate the superiority of the proposed approach in generating contextually accurate event descriptions and mitigating VLM hallucinations. The code will be available at https://github.com/datadrivenwheels/ScVLM
@@ 1 Introduction
In the domain of traffic safety and automatic driving, vision language models (VLMs) have demonstrated strong and robust capabilities in perception, scene understanding, decision-making, and adaptability to novel scenarios [1, 2, 3, 4]. VLMs can proficiently interpret environmental information surrounding the vehicle and possess foundational insights into traffic accidents and potential risk factors [2, 4, 5]. However, despite these advances, challenges still exist in accurately identifying safety-critical events (SCEs), including crashes and near-crashes. Furthermore, understanding the nature of these SCEs, such as conflicts with a lead vehicle, remains elusive. This information is crucial for assessing driving safety.
Figure 1: Example scene understanding result by VideoLLaMA2 (red highlights are the incorrect answers).
Figure 1 illustrates the capabilities of an advanced VLM, VideoLLaMA 2 [6], in understanding SCEs. This model exhibits an excellent understanding of static environmental contexts, including weather conditions and the immediate surroundings. However, its ability to discern dynamic elements crucial for SCE analysis—such as distinguishing between crash and normal driving scenarios or identifying the nature of conflicts (e.g., with a leading or parked vehicle)—is still constrained. These findings underscore the necessity for improved models capable of more effectively interpreting dynamic information in SCE videos.
The limitations observed in VideoLLaMA 2’s performance on SCE analysis can be attributed to two key factors. (1) The rarity of SCEs in real-world scenarios results in insufficient training data for general VLMs to establish connections between SCE videos and corresponding narratives [7, 8]. (2) The scarcity of relevant training examples can lead to hallucinations and the omission of crucial safety characteristics in the model’s interpretations. Additionally, the abstract nature of event types and conflict types poses a significant challenge for VLMs to accurately identify scenarios [9].
This work introduces a novel hybrid approach for generating narratives from driving videos, with a focus on SCEs. The approach combines supervised learning, contrastive learning, and language models to provide accurate and coherent event descriptions. Supervised learning is employed for event type identification (i.e., crashes, tire strikes, near-crashes, normal driving), taking advantage of its effectiveness for task-specific classification. For conflict type identification, contrastive learning is used to capture semantic dependencies between labels and rich textual information. To interpret environmental context, a VLM is utilized to accurately recognize concrete objects within the video. Finally, a Large Language Model (LLM) integrates the outputs from the supervised and contrastive learning components, along with the environmental context, to generate coherent event narratives.
@@
The evaluation of the proposed approach utilized data from the Second Strategic Highway Research Program (SHRP 2) Naturalistic Driving Study (NDS), which is the largest publicly accessible NDS dataset to date, containing over 1 million hours of continuous driving data [10]. The SHRP 2 NDS data includes rich driving information from multiple cameras, kinematic sensors, radar, and GPS. From the continuous driving data, a dedicated project was conducted to identify SCEs and randomly selected normal driving baselines [10], including four distinct event types: crashes, tire strikes, near-crashes, and normal driving baselines. SCEs went through a rigorous data annotation process to extract the nature of the conflict. The annotations provide detailed conflict type labels for SCEs, covering scenarios like conflicts with a lead vehicle, single-vehicle conflicts, and conflicts with a vehicle turning into another’s path in the same direction. This rich dataset is ideal for evaluating the effectiveness of the proposed hybrid approach.
@@ 2 Related Works
VLM for Driving Scene Understanding VLMs combine visual and language processing to interpret driving scenarios and aid decision-making. DriveVLM incorporates reasoning modules for scene description and analysis, addressing spatial reasoning and computational challenges by proposing a hybrid system that combines VLMs with traditional autonomous driving pipelines [3]. DriveScenify utilizes advanced VLMs to generate contextually relevant responses based on driving scene videos, aiming to enhance urban mobility and road safety [11]. Shoman et al. [2] propose a parallel architecture that integrates object detection, tracking, and natural language generation to produce detailed descriptions of traffic events, thereby improving traffic safety through comprehensive event analysis. Jain et al. [12] integrate VLMs with multi-sensor data to enhance the comprehension of traffic dynamics and interactions among road users and infrastructure.
While research on VLMs for general scene understanding in driving contexts has expanded significantly, the specific focus on SCEs, which are vital for improving safety and reliability in autonomous vehicles, remains under-explored. Even if some works mention crashes or traffic accidents [2, 4, 3], they do not explore the intricacies of these events in depth.
Supervised Learning and Contrastive Learning Supervised learning and contrastive learning are two popular approaches for driving video scene classification tasks [13, 14, 15, 16, 17]. Supervised learning relies on one-hot or figure-coded labels to train models [18], while contrastive learning, particularly in a video-text manner, takes advantage of the relationships between different modals of the data to learn useful representations [19]. In supervised learning, state-of-the-art and efficient methods such as SlowFast [20], Swin Transformer [21], and TimeSformer [22] have proven effective for video scene understanding. In contrastive learning, inspired by CLIP [23], notable approaches like X-CLIP [24] and ActionCLIP [25] excel in video understanding, particularly in few-shot tasks. X-CLIP introduced a lightweight cross-frame attention mechanism and proposed a video-adaptive textual prompting scheme to handle video-text datasets [24]. ActionCLIP introduced textual and visual adapters to enhance the model’s ability to process and understand text and video modalities [25].
@@ 3 VLM-based Driving SCE Analysis
Figure 2: The proposed multi-stage approach for generating narrative descriptions of SCEs from driving videos. The process integrates supervised learning for event classification (e.g., crash, near-crash) and contrastive learning for conflict type identification (e.g., conflict with lead vehicle, single vehicle conflict). The VLM extracts visual and environmental information, which is further refined by an LLM to produce a detailed narrative of the SCE.
The proposed approach for generating comprehensive and accurate descriptions of SCEs comprises four distinct stages, as depicted in Figure 2. Initially, a general VLM is employed to extract environmental information from event videos. Subsequently, a supervised learning approach classifies front-view video into four categories: crashes, tire strikes, near-crashes, and normal driving. The third stage utilizes a contrastive learning approach to identify 16 distinct conflict types, such as conflict with a lead vehicle, parked vehicle, and following vehicle. Finally, the framework integrates event classification, conflict type, and environmental context into an LLM to synthesize a comprehensive event description.
@@ 3.1 Supervised Learning for Event Type Classification
Supervised learning for event type classification from video is a 1-of-N vote problem, as illustrated in Figure 3. This type of model takes a video as input and feeds it through a video encoder to generate video representation. The representation is subsequently processed by a classifier to produce prediction scores. The model is optimized by minimizing the cross entropy loss based on the prediction scores. Given an input video $x$ and a label $y$ from a predefined set of labels $Y$, supervised learning approaches typically estimate the model parameter $\theta$ to compute the conditional probability $P(y|x,\theta)$.
@@
The supervised learning approach employs a video encoder $d_{V}$, which extracts representations for video data. Then, the classifier projects the video representations into the space with the dimension of labels to obtain the prediction scores:
$\tilde{y_{et}}=\text{Classifier}[d_{V}(x)]$ (1)
@@
$\begin{split}L=\text{Cross Entropy}[\tilde{y_{et}},y]\end{split}$ (2)
@@ 3.2 Contrastive Learning for Conflict Type Classification
The contrastive learning approach is illustrated in Figure 4. This approach processes a video-text pair as input. The input video is fed into the video encoder to generate video representations. Concurrently, the label text is fed into the text encoder to obtain text representations. The contrastive learning approach computes a similarity score matrix between the video and text representations and is optimized by minimizing the loss between this similarity matrix and the ground-truth video-text pair matrix.
@@
In contrastive learning, the video classification task is redefined as predicting the probability $P[f(x,y)|\theta]$, where $y$ represents the original label texts, $\theta$ refers to model parameters, and $f$ denotes a similarity function. Subsequently, the inference becomes a matching process, with the label texts having the highest similarity score being the outcome:
$\hat{y_{ct}}=\arg\max_{y\in Y}P[f(x,y)|\theta]$ (3)
A contrastive learning approach employs separate encoders $g_{V}$ and $g_{T}$ for videos and label texts within a dual-stream framework. The video encoder $g_{V}$ extracts spatio-and-temporal representations for video data, while the language encoder $g_{T}$ captures representations from label texts. To bring matched video and label representations closer, the similarity score is defined using cosine distances:
@@
$\begin{split}p_{x\to y}(x)=\text{SoftMax}[s(x,y)]\\ p_{y\to x}(y)=\text{SoftMax}[s(y,x)]\end{split}$ (5)
The ground-truth similarity scores are denoted as $q_{x\to y}(x)$ and $q_{y\to x}(y)$, respectively. The negative pair has a similarity of 0, and the positive pair has a similarity of 1. The video-text contrastive loss to be optimized is defined as
@@
A model trained by the contrastive learning approach can carry out inference, as illustrated in Figure 5. When presented with a testing dataset with a label set comprising $M$ labels, the initial step involves extracting the label representations, $[t_{k}],k=1,2,...,M$, using the text encoder, $g_{T}$. Subsequently, for a given testing video, its representation $v$ is obtained through the video encoder, $g_{V}$. The similarity between $v$ and each label representation $t_{k}$ is computed using Equation (4). The label assigned to the video is the one with the highest similarity score with $v$.
@@
3.3 Language Models for Event Narrative Generation
In this study, a VLM is utilized to generate narrative descriptions based on environmental information, such as weather conditions, geographical location, and surrounding context. The process involves the VLM performing inference when provided with a text prompt and the result of SCE detection, enabling accurate event description.
The video representation $r$ for a given input video $x$ is obtained using VLM’s video encoder:
$\begin{split}r=\text{VLM VideoEncoder}(x)\end{split}$ (7)
The representation $r$ is subsequently processed through the Spatial-Temporal Convolution (STC) connector to capture spatial-temporal dynamics. Given a text prompt $P$, along with the predicted event type $\hat{y_{et}}$ and conflict type $\hat{y_{ct}}$ (if applicable), the output response $R$ is generated from an LLM:
$\begin{split}R=\text{LLM}[\text{STC}(r),P,\hat{y_{et}},\hat{y_{ct}}]\end{split}$ (8)
@@ 4 Application and Results
Problem Setup Utilizing the SHRP 2 NDS dataset [10], this study focused on generating accurate narratives from front-view video of SCEs. In the SHRP 2 data set, normal driving segments were captured a few seconds before SCEs within the trip as a reference level. The dataset includes 1,063 crashes, 774 tire strikes, 6,782 near-crashes, and 8,497 normal driving segments. Each event consists of 30-seconds of front-view video. The SCEs are classified into 16 conflict types, as shown in Table 1.
ID Conflict type Count
1 Conflict with a lead vehicle 3165
2 Single vehicle conflict 1441
3 Conflict with vehicle turning into
another vehicle path (same direction) 377
4 Conflict with parked vehicle 173
5 Conflict with vehicle in adjacent lane 1508
6 Conflict with vehicle turning across
another vehicle path (opposite direction) 242
7 Conflict with a following vehicle 181
8 Conflict with vehicle turning into
another vehicle path (opposite direction) 316
9 Conflict with vehicle moving across
another vehicle path (through intersection) 170
10 Conflict with animal 360
11 Conflict with vehicle turning across
another vehicle path (same direction) 65
12 Conflict with merging vehicle 121
13 Conflict with pedal cyclist 64
14 Conflict with pedestrian 163
15 Conflict with obstacle/object in roadway 176
16 Conflict with oncoming traffic 78
17 Unknown 19
Table 1: Count of SCEs by conflict types.
The proposed approach aims to address three tasks: (1) a classification task to distinguish event types, (2) a classification task to differentiate conflict types, and (3) a text generation task to produce event narratives. To the best of the authors’ knowledge, this is the only publicly available driving video dataset with labeled event and conflict types suitable for these three tasks, thereby supporting the evaluation of the proposed approach.
@@ 4.1 SHRP 2 NDS Dataset
Table 2: Comparison of different models in event type classification.
Model & Training set Learning Approach Accuracy Top5 Acc mAP AUC Balanced Acc Macro Precsion Macro F1
X-CLIP (full) Contrastive 0.766 0.951 0.547 0.921 0.493 0.599 0.508
Action CLIP (full) Contrastive 0.748 0.949 0.488 0.907 0.439 0.520 0.458
SlowFast (full) Supervised 0.721 0.928 0.467 0.927 0.437 0.469 0.423
Swin Transformer (full) Supervised 0.719 0.927 0.432 0.889 0.411 0.450 0.420
TimeSformer (full) Supervised 0.713 0.945 0.468 0.920 0.448 0.487 0.459
X-CLIP (5%) Contrastive 0.636 0.847 0.278 0.770 0.232 0.259 0.222
Action CLIP (5%) Contrastive 0.606 0.846 0.239 0.777 0.216 0.244 0.220
SlowFast (5%) Supervised 0.485 0.806 0.148 0.659 0.130 0.156 0.105
Swin Transformer (5%) Supervised 0.545 0.829 0.197 0.752 0.163 0.158 0.153
TimeSformer (5%) Supervised 0.571 0.840 0.205 0.786 0.166 0.177 0.154
Table 3: Comparison of different models in conflict type classification.
The SHRP 2 NDS is the largest naturalistic driving study to date, involving over 3,000 participants and collecting data from vehicles equipped with a comprehensive data recording system [10, 26]. This system captured continuous video footage at 15 FPS from four camera angles, resulting in over a million hours, or 70 million miles, of driving data. SCEs, including crashes, tire strikes, and near-crashes, were identified through kinematic data analysis and video verification [10]. Near-crashes are defined as situations requiring evasive maneuvers to avoid a crash[10], while tire strikes are linked to road departure incidents [27]. The extensive dataset and detailed classification of SCEs, available on the SHRP 2 InSight website [28], provide valuable insights into real-world driving behaviors and safety-critical situations.
@@
Data Pre-processing The time of each SCE was pinpointed using the impact timestamp from the SHRP 2 database and serves as the center point of the event [28]. A temporal window that included 38 video frames (equivalent to 2.5 seconds) both preceding and succeeding the event was extracted, resulting in a 5-second interval of the front-view video. For each SCE, a matched normal driving segment with the same duration as an SCE was randomly selected from the same trip prior to the SCE.
4.2 Classification Task Implementation and Performance